OpenStudy (anonymous):

How would you factor -(3x^2 - 4x + 15)? Explain, please!

OpenStudy (anonymous):

\(-(3x+a)(x+b) \) where \(ab=15\) and \(3b+a=-4\)

OpenStudy (anonymous):

actually this one doesn't factor using integers

OpenStudy (anonymous):

So wait, what. :c

OpenStudy (anonymous):

\( -(3x^2 - 4x + 15)\) does not factor

OpenStudy (anonymous):

The original problem was \[ \frac{ (x^2 + 5) \times 3 - (3x + 2) \times 2x }{ (x^2 + 5)^2 }\]

OpenStudy (anonymous):

So I made the numerator (3x^2 + 15) - (6x^2 +4x)

OpenStudy (anonymous):

\[\frac{ (x^2 + 5) \times 3 - (3x + 2) \times 2x }{ (x^2 + 5)^2 }\]\[=\frac{ 3x^2+15 -6x^2-4x}{ (x^2 + 5)^2 }\]

OpenStudy (zzr0ck3r):

3x^2+15-6x^2-4x = -3x^2-4x+15 = -(3x^2+4x-15)

OpenStudy (anonymous):

combine like terms get \[\frac{-3x^2-4x+15}{(x^2+5)^2}\]

OpenStudy (anonymous):

Which is what I thought I had to factor o:

OpenStudy (zzr0ck3r):

-(3x^2+9x-5x-15) = -(3x(x+3)-5(x+3) = -(3x-5)(x+3)

OpenStudy (anonymous):

\[-3x^2-4x+15=-(3x^2+4x+15)=-(x+3)(3x-5)\]

OpenStudy (anonymous):

it was \(+4x\) not \(-4x\) when you factor out the \(-\)

OpenStudy (anonymous):

Oh! Okay, thank you guys. (: